the gcd of two positive integer is 81 and their lcm is 5103 , find the number

Grade:

3 Answers

anand kapare
9 years ago
i want answer from you
 
anand kapare
9 years ago
find gcd of 15,20,35 . also find x, y, z such that 15x+20y+35z= gcd
Karan Yadav
9 years ago
729 and 567
GCD=81  Therefore numbers are 81a and 81b for a and b are co-prime.
LCM=81*a*b  = 5103=81*63
therefore ab=63 a,b=9,7
so numbers are 81*7=567 and 81*9 =729
 
GCD of 15,20,35 is 5
x=1/3 y=0 z=0